OpenStudy (anonymous):

Please help me!! Write a function rule that represents each sentence. 1. y is 5 less than the product of 4 and x 2. 2.5 more than the quotient of h and 3 is w Write a function rule that represents each situation 1. Pizza the price p of a pizza is $6.95 plus $.95 for each topping t on the pizza

OpenStudy (anonymous):

y is 5 less than the product of 4 and x is means \(=\) 5 less means \(-5\) the product means \(\times \) so \[y=4x-5\]

OpenStudy (anonymous):

can u help me on the other too please?

OpenStudy (anonymous):

quotient of \(h\) and \(3\) means \(\frac{h}{3}\)

OpenStudy (anonymous):

2.5 more means \[\frac{h}{3}+2.5\]

OpenStudy (anonymous):

is \(w\) means \[\frac{h}{3}+2.5=w\]

OpenStudy (anonymous):

I need help on the last one

OpenStudy (skullpatrol):

Hi

OpenStudy (anonymous):

Hey

OpenStudy (skullpatrol):

Pizza the price p of a pizza is $6.95 plus $.95 for each topping t on the pizza.

OpenStudy (anonymous):

Yea i need help with that

OpenStudy (skullpatrol):

What letter are you going to choose for the total price of the pizza?

OpenStudy (anonymous):

x

OpenStudy (skullpatrol):

But, read the problem carefully...

OpenStudy (anonymous):

p

OpenStudy (skullpatrol):

Very good choice :)

OpenStudy (anonymous):

Haha i wasnt thinking at first

OpenStudy (skullpatrol):

You have to keep thinking when you do math.

OpenStudy (anonymous):

I know aha

OpenStudy (skullpatrol):

So, let's write out what you have chosen as Let p = the price of the pizza.

OpenStudy (anonymous):

k

OpenStudy (skullpatrol):

Using that how would you write: "the price p of a pizza is $6.95"

OpenStudy (anonymous):

would it be p=$6.95 ?

OpenStudy (skullpatrol):

Yes, it would be p=$6.95

OpenStudy (anonymous):

Yay!

OpenStudy (skullpatrol):

But that is just the pizza with no toppings, they cost extra, right?

OpenStudy (anonymous):

yea extra $.95

OpenStudy (skullpatrol):

per topping

OpenStudy (skullpatrol):

so it is $$\frac{$.95}{topping}$$

OpenStudy (anonymous):

Yes

OpenStudy (skullpatrol):

What letter are you going to use for the number of toppings on the pizza?

OpenStudy (anonymous):

t

OpenStudy (skullpatrol):

Nice choice :)

OpenStudy (anonymous):

thnx

OpenStudy (skullpatrol):

t can be 1, 2, 3, 4,...

OpenStudy (anonymous):

k

OpenStudy (skullpatrol):